OpenStudy (anonymous):

Cami cycles 7/10 mi each way between her home and school. Victor cycles 3/10 mi less than that each way between his home and school. How much farther does Cami cycle than Victor in 5 days?

OpenStudy (anonymous):

If Cami cycles 7/10 miles each way, she will ride double that each trip \[2(\frac{ 7 }{ 10 }miles)=\frac{ 14 }{ 10 }miles=\frac{ 7 }{ 5 }miles\] Victor cycles 3/10 miles less than Cami does each way, which is \[\frac{ 7 }{ 10 }miles-\frac{ 3 }{ 10 }miles=\frac{ 4 }{ 10 }miles=\frac{ 2 }{ 5 }miles\] Since that's each direction, double that to get how many miles Victor cycles each day \[2(\frac{ 2 }{ 5 }miles)=\frac{ 4 }{ 5 }miles\] So we have that Cami cycles 7/5 miles a day and Victor cycles 4/5 miles a day. If they each ride 5 days, just multiply the distances by 5. So over 5 days, Cami rides a total of: \[5(\frac{ 4 }{ 5 }miles)=\frac{ 20 }{ 5 }miles=4 miles\] And over 5 days, Victor rides a total of: \[5(\frac{ 2 }{ 5 }miles)=\frac{ 10 }{ 5 }miles=2miles\] Subtract the two totals to find how many more miles Cami rode than Victor: \[4 miles - 2 miles = 2 miles\] So you have that Cami rides 2 more miles than Victor over a course of 5 days.